# Magnetic disk winds in protoplanetary disks:

### v2

@@ 3.3 Magnetic disk winds and observational constraints
The disk size and the mass are some of the most fundamental characteristics of a protoplanetary disk and some of the observable population-level differences between disks evolving with and without magnetic winds manifest in these global properties. In this section, we will compare our magnetic disk wind models against those evolving only with gravitoviscous forces, in the context of observational constraints. Here we briefly summarize the complications of observing protoplanetary disks (see Miotello et al. (2023) for a recent review). When measuring disk size, the dust radius is typically measured from surface brightness profile of dust continuum emission in millimeter wavelengths; the most favored for ALMA surveys is Band 6 at 1.3 mm. The radius is defined empirically as the radius which contains a certain fraction of the total disk emission (68%, 90%, or 95%). The gas emissions from a protoplanetary disk can be measured via molecular line traces such as CO isotopologs. For measuring the disk size in gas, the zero moment map of 12CO (J=2–1) is typically utilized, which is optically thick and easier to detect in small amounts at large radii. The gas radius is calculated similar to the methods as described above and it is observed to be larger than the dust radius, typically by a factor of few (Ansdell et al., 2018). As discussed in Section 3.2, the larger gas disk radii can be attributed to optical depth effects as well as dust drift.
@@
With the assumptions of optically thin emission along with a prescription of opacity and disk temperature, the continuum flux can be translated into the dust mass (Hildebrand, 1983), although these assumptions do not always necessarily hold (Dunham et al., 2014; Woitke et al., 2019; Haworth, 2021). Direct measurements of the gas mass, which is expected to contain the majority of the disk mass, are difficult due to the faintness and inaccessibility of H2 line observations. However, the method of measuring gas masses using a tracer such as CO lines has severe limitations arising from the uncertainty in the CO/H ratio in disks (more specifically, gas-phase [CO]/[H2] ratio, see Fig. 3), which is inevitably connected to the disk thermochemistry (Ansdell et al., 2016; Bosman et al., 2021). Traditionally, the dust mass measured from continuum flux is translated into the total disk mass through a multiplicative factor of 100, as indicated by the interstellar gas-to-dust ratio. Predictions from hydrodynamic simulations of PPDs, however, point towards significant local variations from this value due to dust dynamics (Takeuchi & Lin, 2002; Vorobyov & Elbakyan, 2019) Thus, these different approaches of measuring the mass do not agree with each other for PPDs around low mass stars, resulting in significant uncertainties.
@@
With these caveats in mind, we compare our results with the latest compilation of ALMA survey data of Class II protoplanetary disks. The star forming regions considered here are Ophiuchus, Taurus, Lupus and Chamaeleon, which have similar, relatively young ($\sim 0.8-3$ Myr) ages and form a near complete sample (Ansdell et al., 2016, 2018; Manara et al., 2023a; Miotello et al., 2023). Since not all data is available for each of the star forming regions, Table 5 summarizes the data used for comparison with the simulated models. The theoretical masses and radii resulting from simulations are not directly observable, which motivates us to make synthetic observations with ProDiMo as elaborated in Section 3.2. Here we compare the synthetic observations in terms of radii $R_{\rm{dust,Band6}}$ and $R_{\rm gas,^{12}{CO}}$, both calculated at 90% flux threshold. The dust content of the disk $M_{\rm dust,Band6}$ is also compared with the survey data, while for the gas component, we use the line fluxes of ${\rm{}^{12}{CO}}$ and ${\rm{C^{13}O}}$ at a standard distance of 150 pc.
@@
Figure 4: Evolution of the disk size – theoretical centrifugal disk radii (top panel) and radii obtained from ProDiMo synthetic observations (bottom two panels). The gas and dust radii for the models are calculated via synthetic 12CO and Band 6 continuum emissions with 90% flux threshold. The vertical, dashed gray lines mark the average Class 0/I and Class I/II boundaries. The shaded regions on the right depict half-violin plots of gas (red) and dust (red) radii of the Class II disks, as observed in a subset of ALMA surveys (see Table 5).
@@
In Fig. 4 we show the temporal evolution of the centrifugal radius as well as synthetic gas and dust radii for the models. The vertical dotted lines in these plots indicate the average location of the Class 0/I and Class I/II boundaries, as defined in Section 3.1. The top panel of this figure shows the theoretical size evolution of the disks in terms of $R_{\rm cf}$, which is smoothed with a box-car average with a box-width of 0.05 Myr, in order to get rid of stochastic variations caused by disk size pulsation inherit to young stellar objects with variable protostellar accretion (Vorobyov et al., 2020a). The significantly smaller sizes of low viscosity disks with winds in WI-3 and WI-4 are apparent in these plots. Even in the early Class 0 and I stages, these disks are much smaller than their gravitoviscous counterparts. As also seen in Fig. 1, the disks in models GV-4, WI-2 and WI-3a cluster around 150-250 au, while GV-2 produces a large viscously spreading disk. The exact values of disk radii as well as some of the global disk properties at 1 Myr are listed in Table 4 for comparison.
@@
The second and third panels of Fig. 4 show synthetic measurements of gas and dust radii, respectively. As elaborated in Section 3.2, the gas radius is approximated from 12CO(J=2-1) line emissions, while the dust radius is estimated from simulated 1.3 mm continuum observations, each with a 90% flux threshold. The synthetic observations obtained via ProDiMo are performed after the embedded Class 0 and Class I stages are over, when the disk may be visible through the nascent cloud code. Starting at 0.2 Myr, a total of five ProDiMo calculations are performed per model, which are spaced uniformly in time, so that the observational trends can be followed as a PPD evolves through the Class II stage. The shaded regions near the right axis of these panels depict half-violin plots of observed gas and dust radius data. The gas radii pertain to Lupus cluster only (Ansdell et al., 2018), while the dust radii are obtained from all four aforementioned star forming regions (Hendler et al., 2020; Manara et al., 2023a). These violin plots show the probability distribution functions for the data, which are terminated at the extreme values denoted by dashes, with the central dash showing the median value.
@@
Figure 5: Evolution of the disk mass – theoretical mass within the centrifugal region (top panel), theoretical dust mass within the centrifugal disk, neglecting the innermost 5 au region (second panel), dust mass obtained from ProDiMo synthetic observations (third panel), and synthetic CO line flux as a proxy for the disk gas content (last two panels). The dust masses are calculated from continuum Band 6 fluxes, while the 13CO and C18O line fluxes are calculated at 150 pc. The vertical, dashed gray lines mark the average Class 0/I and Class I/II boundaries. The shaded regions on the right depict half-violin plots of dust masses (red) and line fluxes (blue) in Class II disks, as observed in a subset of ALMA surveys. The vertical, dashed gray lines mark the average Class 0/I and Class I/II boundaries.
@@
Considering the evolution of dust radius in the third panel of Fig. 4, the GV disks show a monotonic increase. With the inclusion of winds, the dust disk radius decreases as a disk enters Class II stage and afterwards, it may increase depending on the model. Recent survey of Orion molecular clouds indicates a similar decrease in the mean dust disk radius with evolution, which supports the scenario of disk winds (Tobin et al., 2020). The impact of dust evolution on the disk radius in the presence of winds can be summarized as follows. The dust size in the outer regions of a PPD is typically drift limited, while in the inner regions the dust may grow to reach the fragmentation barrier (Birnstiel et al., 2016). As a consequence of the dust partitioning, in the winds in the outer disk evacuate a relatively larger amount of (small) dust per unit mass. The remaining grown dust is subject to efficient radial drift towards increasing gas pressure in the low $\alpha$ environment, since a lower $\alpha$ increases dust fragmentation barrier, dust maximum size, and the Stokes number. PPDs subject to photoevaporative winds cause a reduction of dust reservoir at large radii by a similar mechanism (Sellek et al., 2020). The bulk of the disk is thus depleted of small dust and grown dust drifts to the inner disk regions where it gets trapped in the gaseous rings, with a large dust-to-gas ratio. As the result, the dust disk size in WI-3 and WI-4 shrinks to only a few au in Band 6 continuum flux. With the adjust wind model WI-3a, the effects of wind are attenuated by increasing $C_{\beta}$ parameter by a factor of 5. This modulates the inwards transport of gas and consequently the dust drift, so that the continuum disk in synthetic observations are congruent with observations even at low $\alpha_{\rm MRI}=10^{-3}$. The dust dynamics with disk winds will be studied in detail in a subsequent paper.
@@
Fig. 5 shows the evolution of disk mass for the models. The true disk gas mass in the model ($M_{\rm disk}$) is plotted in the first panel, which is calculated by summation over the extent of the centrifugally supported disk. All disks achieve the maximum disk mass at the beginning of the Class I stage and afterwards in the Class II stage, the masses decrease monotonically at different rates. As a consequence, the disk mass is typically a substantial fraction of the stellar mass during the Class I stage and the disk is prone to gravitational instability (see second column of Fig. 1). In addition to being large in size, the GV models typically produce more massive disks, especially at low viscosity. With a low value of $\alpha_{\rm MRI}=10^{-4}$ in model GV-4, the transport of angular momentum due to viscous torques is diminished, which over time results in the accumulation of material and a more massive disk than GV-2. Both models WI-3 and WI-4 produce disks with masses in the vicinity of $10^{-4}M_{\odot}$ or 0.1 $M_{\rm J}$. Thus, these disks do not have sufficient mass to produce giant planets, which accrete gas over timescale of several Myr in the core accretion scenario (e.g., Pollack et al., 1996). Models WI-2 and WI-3a produce disks that are of the order of 0.01 $M_{\odot}$ and are most consistent with our understanding of PPD masses in the T Tauri phase.
@@
A key trend observed in the first panel of Fig. 5 is that the lower viscosity disks with gravitoviscous evolution are more massive. However, the trend is reversed with the inclusion of disk winds, i.e., lower viscosity disks get progressively less massive over time. This can be explained with the strong dependence of disk wind mass loss and torque on the gas surface density. The low viscosity creates both mass and angular momentum transport bottleneck, and forces the disk to have a larger $\Sigma_{\rm g}$, which in turn produces strong winds with more torque as well as mass loss. The net effect is that the winds remain strong, until the surface density is reduced, which ultimately results in a less massive disk. This should increase the mass of the central star in wind models relatively rapidly, which is seen in Table 4.
@@
The second panel of Fig. 5 shows the theoretical dust mass of the disk, calculated as the sum of both small and grown dust components. This is done with an exclusion of the innermost 5 au of the disk, in order to avoid the region where the disk shows formation of dusty rings (Kadam et al., 2019, 2022). These rings form pressure maxima and tend to foster dust growth up to meter-sized boulders, which are not observable and contain significant amount of dust mass (Vorobyov et al., 2024). Additionally, the innermost region is typically optically thick, and does not reveal its true mass in observed disks. Thus, removing the innermost region from our theoretical dust mass estimates seems to be the best compromise, for comparison with the synthetic observations. For models WI-3 and WI-4, the dust radius collapses after about 0.5 Myr (last panel of Fig. 4), implying that the calculated dust masses outside of 5 au are correspondingly small and not representative of the dust content of the disk. However, for the rest of the models, the theoretical dust masses serve as a good proxy for the bulk of the observable dust mass in the disk.
@@
The third panel of Fig. 5 shows the synthetically derived dust masses of the simulated disks in comparison with observations. The red violin plot in this panel shows the observed distribution of dust masses as calculated from ALMA Band 6 continuum data, where all four nearby star forming regions are considered (Miotello et al., 2023; Manara et al., 2023a). The abrupt cutoff at the lower end of the observed kernel density distribution is caused by the detection limit of the surveys. Similar to the ALMA surveys, we assume that the dust emitting at sub-millimeter wavelengths is optically thin and it emits isothermally at a characteristic temperature of $T_{\rm d}=20$ K (Hildebrand, 1983; Andrews & Williams, 2005; Manara et al., 2023a). This allows us to estimate the disk mass from the Band 6 (1.3 mm) flux ($F_{\nu}$) as
@@
As mentioned earlier, gas mass of a PPD may be inferred by multiplying its dust mass by a factor of 100. If we calculate the gas masses this way, the estimates will follow the exact same trends as the third panel of Fig. 5, shifted up by the same factor. As estimating disk gas mass comes with significant uncertainties, for the most transparent and direct comparison against the observations, the last two panels of Fig. 5 show the 13CO and C18O line fluxes at 150 au. The observed ranges are obtained from ALMA survey data for Lupus and Chamaeleon star forming regions (Ansdell et al., 2016; Long et al., 2017). In general, the line fluxes from the models are larger than the observed distributions and there may be several contributing factors. Firstly, we note that in ProDiMo, CO isotope-selective processes are not considered and the abundance of CO isotopologues is inferred with fixed ratios of $[^{13}{\rm C}]/[^{12}{\rm C}]=0.0140$ and $[^{18}{\rm O}]/[^{16}{\rm O}]=0.0020$. The less abundant CO isotopologues are thought to become optically thick much deeper in a PPD, thereby tracing the entire thickness of bulk of the disk. However, the increase in line flux with time for the higher mass models suggests that both of these traces are optically thick and therefore follow the increase in disk size with gravitoviscous spread. The simulated disks evolve only up to 1 Myr, while the considered star forming regions may be up 3 Myr old. As a PPD looses its mass with age, this would decrease the observed CO fluxes with respect to synthetic observations. The ProDiMo fluxes are calculated for a face-on disk with zero inclination, which may contribute to a higher CO flux as compared with the observed sample of disks. Finally, we note that CO observations imply a bulk gas-to-dust ratio between 1-10 for most disks, which is much lower than canonical MMSN or ISM value of 100 (Miotello et al., 2017). This discrepancy is attributed to carbon depletion in PPDs, with processes such as CO photodissociation in the upper layers and freeze-out near the midplane. If the CO emissions are optically thick and dust emission optically thin, this will also result in an underestimation of the measured bulk gas-to-dust ratio as compared to the actual value. The CO fluxes demonstrate that the adjusted disk wind model does not fit all observables simultaneously. Nonetheless, with the inclusion of disk winds, the simulated line fluxes decrease and move in the correct direction with respect to an agreement with the observed distributions.
@@
Figure 6: Comparison of the synthetic observations obtained from the disk models with the ALMA survey data with respect to the spectral indices (first panel) and dust-gas radii (second panel). The colored lines show evolutionary sequence of the models, starting at 0.2 Myr and ending at 1 Myr, with the latter point marked with a diamond. The blue scatter plots show observed quantities in Lupus cluster (Ansdell et al., 2018).
@@
We now test the disk evolution predicted by our models against observed spectral indices and the ratio of dust and gas disk sizes. Fig. 6 compares the disk models with ALMA survey data of Class II PPDs in Lupus cluster (Ansdell et al., 2018). The top panel in Fig. 6 shows the total continuum flux versus far infrared spectral index between ALMA Band 6 (1.3 mm) and 7 (0.9 mm). The latter is the SED slope defined as
@@
where $F_{\nu}$ is the continuum flux at the corresponding ALMA band. The $\alpha_{\rm mm}=2$ line is characteristic of large particles represented by blackbody emission, since $F_{\nu}\propto\nu^{2}$. The ISM value of 3.7 corresponding to small particles is also shown in the plot. Assuming optically thin emission, the flux in sub-mm continuum is a direct proxy of the total dust content in the disk. The spectral index probes characteristic grain size and hence often called the “growth index”. Although there is a higher degree of uncertainty, as additional factors such as optical depth and dust scattering complicate the picture (Draine, 2006). ProDiMo calculates the fluxes at ALMA Bands at a distance of 150 pc and the spectral index is calculated from Eq. (30). We note that the observations have large error bars in spectral index, which are not shown in the figure. A disk starts at 0.2 Myr, when it is fully formed and typically evolves diagonally from lower right to upper left this phase space (diamond in the figure represents the end state at 1 Myr). Similar trends in the disk evolution have been previously shown in several studies (Birnstiel et al., 2010; Pinilla et al., 2012; Akimkin et al., 2020). The decrease in Band 6 flux indicates a decrease in dust mass or, in the case of optically thick emissions, an area effect due to inward dust drift. The evolution of spectral index is more complicated; the initial spectral index near 2 is a result of initial large optical depths and its value increases when the grain size approaches $\approx\lambda/2\pi$ of the corresponding wavelength. The disks with large viscous $\alpha$ tend to show both larger flux and spectral index. The values of both spectral index and continuum flux obtained from the models are consistent with those in the ALMA surveys throughout the first 1 Myr evolution. The $\alpha_{\rm mm}$ for WI-3 and WI-4 models first increases and then decreases with time. This decrease at later evolutionary stages is very likely a signature of dust clearing process, as all big grains drift inward to the pressure maxima (see Section 3.2). The $\alpha_{\rm mm}$ is essentially disk-averaged spectral index and corresponds to the most luminous parts of the disk. As most of the dust mass in these low $\alpha$ models with winds eventually accumulates in the inner rings, the associated increase in the average optical depth may be responsible for the re-decrease in $\alpha_{\rm mm}$ at later times.

### v3

@@ 3.3 Magnetic disk winds and observational constraints
The disk size and the mass are some of the most fundamental characteristics of a protoplanetary disk and some of the observable population-level differences between disks evolving with and without magnetic winds manifest in these global properties. In this section we compare our magnetic disk wind models against those evolving only with gravitoviscous forces, in the context of observational constraints. Here we briefly summarize the complications of observing protoplanetary disks (see Miotello et al. (2023) for a recent review). When measuring disk size, the dust radius is typically measured from surface brightness profile of dust continuum emission in millimeter wavelengths; the most favored for ALMA surveys is Band 6 at 1.3 mm. The radius is defined empirically as the radius which contains a certain fraction of the total disk emission (68%, 90%, or 95%). The gas emissions from a protoplanetary disk can be measured via molecular line traces such as CO isotopologs. For measuring the disk size in gas, the zero moment map of 12CO (J=2–1) is typically utilized, which is optically thick and easier to detect in small amounts at large radii. The gas radius is calculated similar to the methods as described above and it is observed to be larger than the dust radius, typically by a factor of a few (Ansdell et al., 2018). As discussed in Sect. 3.2, the larger gas disk radii can be attributed to optical depth effects as well as dust drift.
@@
With the assumptions of optically thin emission along with a prescription of opacity and disk temperature, the continuum flux can be translated into the dust mass (Hildebrand, 1983), although these assumptions do not always necessarily hold (Dunham et al., 2014; Woitke et al., 2019; Haworth, 2021). Direct measurements of the gas mass, which is expected to contain the majority of the disk mass, are difficult due to the faintness and inaccessibility of H2 line observations. However, the method of measuring gas masses using a tracer such as CO lines has severe limitations arising from the uncertainty in the CO/H ratio in disks (more specifically, gas-phase [CO]/[H2] ratio, see Fig. 3), which is inevitably connected to the disk thermochemistry (Ansdell et al., 2016; Bosman et al., 2021). Traditionally, the dust mass measured from continuum flux is translated into the total disk mass through a multiplicative factor of 100, as indicated by the interstellar gas-to-dust ratio. Predictions from hydrodynamic simulations of PPDs, however, point toward significant local variations from this value due to dust dynamics (Takeuchi & Lin, 2002; Vorobyov & Elbakyan, 2019) Thus, these different approaches of measuring the mass do not agree with each other for PPDs around low mass stars, resulting in significant uncertainties.
@@
With these caveats in mind, we compare our results with the latest compilation of ALMA survey data of Class II protoplanetary disks. The star forming regions considered here are Ophiuchus, Taurus, Lupus and Chamaeleon, which have similar, relatively young ($\sim 0.8-3$ Myr) ages and form a near complete sample (Ansdell et al., 2016, 2018; Manara et al., 2023a; Miotello et al., 2023). Since not all data is available for each of the star forming regions, Table 5 summarizes the data used for comparison with the simulated models. The theoretical masses and radii resulting from simulations are not directly observable, which motivates us to make synthetic observations with ProDiMo as elaborated in Sect. 3.2. Here we compare the synthetic observations in terms of radii $R_{\rm{dust,Band6}}$ and $R_{\rm gas,^{12}{CO}}$, both calculated at 90% flux threshold. The dust content of the disk $M_{\rm dust,Band6}$ is also compared with the survey data, while for the gas component, we use the line fluxes of ${\rm{}^{12}{CO}}$ and ${\rm{C^{13}O}}$ at a standard distance of 150 pc.
@@
Figure 4: Evolution of the disk size. Top panel: Theoretical centrifugal disk radii. Bottom two panels: Radii obtained from ProDiMo synthetic observations . The gas and dust radii for the models are calculated via synthetic 12CO and Band 6 continuum emissions with 90% flux threshold. The vertical, dashed gray lines mark the average Class 0/I and Class I/II boundaries. The shaded regions on the right depict half-violin plots of gas (red) and dust (red) radii of the Class II disks, as observed in a subset of ALMA surveys (see Table 5).
@@
In Fig. 4 we show the temporal evolution of the centrifugal radius as well as synthetic gas and dust radii for the models. The vertical dotted lines in these plots indicate the average location of the Class 0/I and Class I/II boundaries, as defined in Sect. 3.1. The top panel of this figure shows the theoretical size evolution of the disks in terms of $R_{\rm cf}$, which is smoothed with a box-car average with a box-width of 0.05 Myr, in order to get rid of stochastic variations caused by disk size pulsation inherit to young stellar objects with variable protostellar accretion (Vorobyov et al., 2020a). The significantly smaller sizes of low viscosity disks with winds in WI-3 and WI-4 are apparent in these plots. Even in the early Class 0 and I stages, these disks are much smaller than their gravitoviscous counterparts. As also seen in Fig. 1, the disks in models GV-4, WI-2 and WI-3a cluster around 150-250 au, while GV-2 produces a large viscously spreading disk. The exact values of disk radii as well as some of the global disk properties at one Myr are listed in Table 4 for comparison.
@@
The second and third panels of Fig. 4 show synthetic measurements of gas and dust radii, respectively. As elaborated in Sect. 3.2, the gas radius is approximated from 12CO(J=2-1) line emissions, while the dust radius is estimated from simulated 1.3 mm continuum observations, each with a 90% flux threshold. The synthetic observations obtained via ProDiMo are performed after the embedded Class 0 and Class I stages are over, when the disk may be visible through the nascent cloud code. Starting at 0.2 Myr, a total of five ProDiMo calculations are performed per model, which are spaced uniformly in time so that the observational trends can be followed as a PPD evolves through the Class II stage. The shaded regions near the right axis of these panels depict half-violin plots of observed gas and dust radius data. The gas radii pertain to Lupus cluster only (Ansdell et al., 2018), while the dust radii are obtained from all four aforementioned star forming regions (Hendler et al., 2020; Manara et al., 2023a). These violin plots show the probability distribution functions for the data, which are terminated at the extreme values denoted by dashes, with the central dash showing the median value.
@@
Figure 5: Evolution of the disk mass. Top panel: Theoretical mass within the centrifugal region. Second panel: Theoretical dust mass within the centrifugal disk, neglecting the innermost five au region. Third panel: Dust mass obtained from ProDiMo synthetic observations. Last two panels: Synthetic CO line flux as a proxy for the disk gas content. The dust masses are calculated from continuum Band 6 fluxes, while the 13CO and C18O line fluxes are calculated at 150 pc. The vertical, dashed gray lines mark the average Class 0/I and Class I/II boundaries. The shaded regions on the right depict half-violin plots of dust masses (red) and line fluxes (blue) in Class II disks, as observed in a subset of ALMA surveys. The vertical, dashed gray lines mark the average Class 0/I and Class I/II boundaries.
@@
Considering the evolution of dust radius in the third panel of Fig. 4, the GV disks show a monotonic increase. With the inclusion of winds, the dust disk radius decreases as a disk enters Class II stage and afterward, it may increase depending on the model. Recent survey of Orion molecular clouds indicates a similar decrease in the mean dust disk radius with evolution, which supports the scenario of disk winds (Tobin et al., 2020). The impact of dust evolution on the disk radius in the presence of winds can be summarized as follows. The dust size in the outer regions of a PPD is typically drift limited, while in the inner regions the dust may grow to reach the fragmentation barrier (Birnstiel et al., 2016). As a consequence of the dust partitioning, in the winds in the outer disk evacuate a relatively larger amount of (small) dust per unit mass. The remaining grown dust is subject to efficient radial drift toward increasing gas pressure in the low $\alpha$ environment, since a lower $\alpha$ increases dust fragmentation barrier, dust maximum size, and the Stokes number. PPDs subject to photoevaporative winds cause a reduction of dust reservoir at large radii by a similar mechanism (Sellek et al., 2020). The bulk of the disk is thus depleted of small dust and grown dust drifts to the inner disk regions where it gets trapped in the gaseous rings, with a large dust-to-gas ratio. As the result, the dust disk size in WI-3 and WI-4 shrinks to only a few au in Band 6 continuum flux. With the adjust wind model WI-3a, the effects of wind are attenuated by increasing $C_{\beta}$ parameter by a factor of 5. This modulates the inward transport of gas and consequently the dust drift so that the continuum disk in synthetic observations are congruent with observations even at low $\alpha_{\rm MRI}=10^{-3}$. The dust dynamics with disk winds will be studied in detail in a subsequent paper.
@@
Figure 5 shows the evolution of disk mass for the models. The true disk gas mass in the model ($M_{\rm disk}$) is plotted in the first panel, which is calculated by summation over the extent of the centrifugally supported disk. All disks achieve the maximum disk mass at the beginning of the Class I stage and afterward in the Class II stage, the masses decrease monotonically at different rates. As a consequence, the disk mass is typically a substantial fraction of the stellar mass during the Class I stage and the disk is prone to gravitational instability (see second column of Fig. 1). In addition to being large in size, the GV models typically produce more massive disks, especially at low viscosity. With a low value of $\alpha_{\rm MRI}=10^{-4}$ in model GV-4, the transport of angular momentum due to viscous torques is diminished, which over time results in the accumulation of material and a more massive disk than GV-2. Both models WI-3 and WI-4 produce disks with masses in the vicinity of $10^{-4}M_{\odot}$ or 0.1 $M_{\rm J}$. Thus, these disks do not have sufficient mass to produce giant planets, which accrete gas over timescale of several million years in the core accretion scenario (e.g., Pollack et al., 1996). However, we note that the planet formation possibly begins before Class II stage and at that time, the mass available for this process is much more than typical assumptions of MMSN disk (Tychoniec et al., 2020; Schib et al., 2021). Models WI-2 and WI-3a produce disks that are of the order of 0.01 $M_{\odot}$ and are most consistent with our understanding of PPD masses in the T Tauri phase.
@@
A key trend observed in the first panel of Fig. 5 is that the lower viscosity disks with gravitoviscous evolution are more massive. However, the trend is reversed with the inclusion of disk winds; that is, lower viscosity disks become progressively less massive over time. This can be explained with the strong dependence of disk wind mass loss and torque on the gas surface density. The low viscosity creates both mass and angular momentum transport bottleneck, and forces the disk to have a larger $\Sigma_{\rm g}$, which in turn produces strong winds with more torque as well as mass loss. The net effect is that the winds remain strong, until the surface density is reduced, which ultimately results in a less massive disk. This should increase the mass of the central star in wind models relatively rapidly, which is seen in Table 4.
@@
The second panel of Fig. 5 shows the theoretical dust mass of the disk, calculated as the sum of both small and grown dust components. This is done with an exclusion of the innermost five au of the disk, in order to avoid the region where the disk shows formation of dusty rings (Kadam et al., 2019, 2022). These rings form pressure maxima and tend to foster dust growth up to meter-sized boulders, which are not observable and contain significant amount of dust mass (Vorobyov et al., 2024). Additionally, the innermost region is typically optically thick, and does not reveal its true mass in observed disks. Thus, removing the innermost region from our theoretical dust mass estimates seems to be the best compromise, for comparison with the synthetic observations. For models WI-3 and WI-4, the dust radius collapses after about 0.5 Myr (last panel of Fig. 4), implying that the calculated dust masses outside of five au are correspondingly small and not representative of the dust content of the disk. However, for the rest of the models, the theoretical dust masses serve as a good proxy for the bulk of the observable dust mass in the disk.
@@
The third panel of Fig. 5 shows the synthetically derived dust masses of the simulated disks in comparison with observations. The red violin plot in this panel shows the observed distribution of dust masses as calculated from ALMA Band 6 continuum data, where all four nearby star forming regions are considered (Miotello et al., 2023; Manara et al., 2023a). The abrupt cutoff at the lower end of the observed kernel density distribution is caused by the detection limit of the surveys. Similar to the ALMA surveys, we assume that the dust emitting at submillimeter wavelengths is optically thin and it emits isothermally at a characteristic temperature of $T_{\rm d}=20$ K (Hildebrand, 1983; Andrews & Williams, 2005; Manara et al., 2023a). This allows us to estimate the disk mass from the Band 6 (1.3 mm) flux ($F_{\nu}$) as
@@
As mentioned earlier, gas mass of a PPD may be inferred by multiplying its dust mass by a factor of 100. If we calculate the gas masses this way, the estimates will follow the exact same trends as the third panel of Fig. 5, shifted up by the same factor. As estimating disk gas mass comes with significant uncertainties, for the most transparent and direct comparison against the observations, the last two panels of Fig. 5 show the 13CO and C18O line fluxes at 150 au. The observed ranges are obtained from ALMA survey data for Lupus and Chamaeleon star forming regions (Ansdell et al., 2016; Long et al., 2017). In general, the line fluxes from the models are larger than the observed distributions and there may be several contributing factors. Firstly, we note that in ProDiMo, CO isotope-selective processes are not considered and the abundance of CO isotopologues is inferred with fixed ratios of $[^{13}{\rm C}]/[^{12}{\rm C}]=0.0140$ and $[^{18}{\rm O}]/[^{16}{\rm O}]=0.0020$. The less abundant CO isotopologues are thought to become optically thick much deeper in a PPD, thereby tracing the entire thickness of bulk of the disk. However, the increase in line flux with time for the higher mass models suggests that both of these traces are optically thick and therefore follow the increase in disk size with gravitoviscous spread. The simulated disks evolve only up to one Myr, while the considered star forming regions may be up three Myr old. As a PPD looses its mass with age, this would decrease the observed CO fluxes with respect to synthetic observations. The ProDiMo fluxes are calculated for a face-on disk with zero inclination, which may contribute to a higher CO flux as compared with the observed sample of disks. Finally, we note that CO observations imply a bulk gas-to-dust ratio between 1-10 for most disks, which is much lower than canonical MMSN or ISM value of 100 (Miotello et al., 2017). This discrepancy is attributed to carbon depletion in PPDs, with processes such as CO photodissociation in the upper layers and freeze-out near the midplane. If the CO emissions are optically thick and dust emission optically thin, this will also result in an underestimation of the measured bulk gas-to-dust ratio as compared to the actual value. The CO fluxes demonstrate that the adjusted disk wind model does not fit all observables simultaneously. Nonetheless, with the inclusion of disk winds, the simulated line fluxes decrease and move in the correct direction with respect to an agreement with the observed distributions.
@@
Figure 6: Comparison of the synthetic observations obtained from the disk models with the ALMA survey data with respect to the spectral indices (first panel) and dust-gas radii (second panel). The colored lines show evolutionary sequence of the models, starting at 0.2 Myr and ending at one Myr, with the latter point marked with a diamond. The blue scatter plots show observed quantities in Lupus cluster (Ansdell et al., 2018).
@@
We now test the disk evolution predicted by our models against observed spectral indices and the ratio of dust and gas disk sizes. Figure 6 compares the disk models with ALMA survey data of Class II PPDs in Lupus cluster (Ansdell et al., 2018). The top panel in Fig. 6 shows the total continuum flux versus far infrared spectral index between ALMA Band 6 (1.3 mm) and 7 (0.9 mm). The latter is the SED slope defined as
@@
where $F_{\nu}$ is the continuum flux at the corresponding ALMA band. The $\alpha_{\rm mm}=2$ line is characteristic of large particles represented by blackbody emission, since $F_{\nu}\propto\nu^{2}$. The ISM value of 3.7 corresponding to small particles is also shown in the plot. Assuming optically thin emission, the flux in submillimeter continuum is a direct proxy of the total dust content in the disk. The spectral index probes the characteristic grain size and is hence often called the “growth index.” Although there is a higher degree of uncertainty, as additional factors such as optical depth and dust scattering complicate the picture (Draine, 2006). ProDiMo calculates the fluxes at ALMA Bands at a distance of 150 pc and the spectral index is calculated from Eq. (30). We note that the observations have large error bars in spectral index, which are not shown in the figure. A disk starts at 0.2 Myr, when it is fully formed and typically evolves diagonally from lower right to upper left this phase space (diamond in the figure represents the end state at one Myr). Similar trends in the disk evolution have been previously shown in several studies (Birnstiel et al., 2010; Pinilla et al., 2012; Akimkin et al., 2020). The decrease in Band 6 flux indicates a decrease in dust mass or, in the case of optically thick emissions, an area effect due to inward dust drift. The evolution of spectral index is more complicated; the initial spectral index near 2 is a result of initial large optical depths and its value increases when the grain size approaches $\approx\lambda/2\pi$ of the corresponding wavelength. The disks with large viscous $\alpha$ tend to show both larger flux and spectral index. The values of both spectral index and continuum flux obtained from the models are consistent with those in the ALMA surveys throughout the first Myr of evolution. The $\alpha_{\rm mm}$ for WI-3 and WI-4 models first increases and then decreases with time. This decrease at later evolutionary stages is very likely a signature of dust clearing process, as all big grains drift inward to the pressure maxima (see Sect. 3.2). The $\alpha_{\rm mm}$ is essentially disk-averaged spectral index and corresponds to the most luminous parts of the disk. As most of the dust mass in these low $\alpha$ models with winds eventually accumulates in the inner rings, the associated increase in the average optical depth may be responsible for the re-decrease in $\alpha_{\rm mm}$ at later times.